OpenStudy (anonymous):

find the first five terms of the given sequence a1=2,an+1=an-7 a. 2,-5,-12,-19,-26 b. 2,-26,-19,-12,-5 c. 2,-9,-16,-23,-30 d. 2,-5,-4,-3,-2

OpenStudy (anonymous):

this is a lot easier than the last one! you can just about do it in your head

OpenStudy (anonymous):

\[a_1=2\] \[a_2=2-7=-5\] \[a_3=-5-7=-12\]...